OpenStudy (anonymous):

27=8(x+5)+5x solve for x I need help please!!!

sammixboo (sammixboo):

Okie dokie =) So does your problem look like this? \(27=8(x+5)+5x\) or \(27=8(x+5)+5^x\)

OpenStudy (anonymous):

@sammixboo the first one

OpenStudy (anonymous):

27=8(x+5)+5x 27=8x+40+5x 27-40=8x+5x -13=13x x=-13/13 x=-1

sammixboo (sammixboo):

OK, so we have 27=8(x+5)+5x First we need to distribute 8(x+5) by multiplying 8 by x and 8 by 5 to get 27=8x + 40 + 5x

sammixboo (sammixboo):

If you want me to explain in a more detailed way let me know @mari_2495

OpenStudy (anonymous):

@Angu thank you can you help me with this one 8v+2(v-8)=4

OpenStudy (anonymous):

follow the same procedure..

sammixboo (sammixboo):

Well, you start off 8v+2(v-8)=4 by distributing the 2 to get 8v + 2v - 16=4 Now you have to simplify by adding 8v + 2v Do you know what 8v + 2v equals?

OpenStudy (anonymous):

@Angu 8v+2v-16=4 10v-16=4 am I doing it right?'ugh I need help im trying

sammixboo (sammixboo):

Right!

sammixboo (sammixboo):

Now do you know what to do next?

OpenStudy (anonymous):

ya.. continue

OpenStudy (anonymous):

@sammixboo no I needhelp

sammixboo (sammixboo):

OK, so all you need to do is ADD 16 to both sides of the equation 10v - 16 + 16 = 4 + 16 10v = 4 + 16 What is 4 + 16?

OpenStudy (anonymous):

20 @sammixboo

sammixboo (sammixboo):

Right, so 10v = 20 Do you know what to do next?

OpenStudy (anonymous):

well do I divide 10

sammixboo (sammixboo):

Yup. 10v / 10 = 20/10 v = 20/10

OpenStudy (anonymous):

@sammixboo so its 2

sammixboo (sammixboo):

Right! v = 2

OpenStudy (anonymous):

@sammixboo thank you :) and can you help me with another different problem please

sammixboo (sammixboo):

One other thing \(\huge \rlap{\bbox[2.9pt, ,border:6.8px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{Welcome}}}}\hspace{127pt}\huge \rlap{\bbox[3.3pt, ,border:7px solid #2aacd7 ]{\Huge\color{blueviolet}{\rm \fbox{to}}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \color{blueviolet}{\heartsuit^{\LARGE \heartsuit^{\Large \heartsuit}}}\) \(\huge \huge \rlap{\bbox[5pt, ,border:8px solid #2aacd7 ]{\Huge\color{#84c73c}{\rm \fbox{OpenStudy}}}}\)

sammixboo (sammixboo):

Yeah sure! Make sure when you post your questions, post them on different posts so we don't spam other users and to make things easier =)

sammixboo (sammixboo):

You have to close this post first to post another question

OpenStudy (anonymous):

okay